Solve the system of equations by addition method
4x+2y=4
8x+4y=0

Answer by unlockmath(1688)

Hello,
To solve this, let's multiply the first equation by -2.
(-2)4x+(-2)2y=(-2)4 that gives us:
-8x-4y=-8 Now we can add this to the second equation.
8x+4y=0 Notice the y and x cancels out leaving us with:
0=-8 This obviously doesn't make sense so there is NO SOLUTION to this problem.
Check out my book at www.math-unlock.com
RJ Toftness